### v1

@@ Results
Medical concept annotation
Figure 2: Visualization of zero-shot medical concept annotation. For each example, the focused regions are highlighted based on the gradient of image-concept similarities, where the predicted probabilities of concepts are also presented.
Figure 3: Concept-based data analysis on datasets of various modalities. We calculate the concept presence difference to obtain which concepts are more likely to appear in which image set using the zero-shot concept annotation capability of our model.
Figure 4: Examples of concept-class association. We present the top three relevant concepts for selected diseases, ranked by their weights in the concept linear layer, paired with an example image of the specific disease from each considered dataset.
Figure 5: Disease-level concept inspection. For a specific disease, disease-level concept inspection conducts concept-disease association analysis from multiple datasets of different institutions to get more accurate and consistent concept-based explanations.
Trustworthiness is essential for AI models to be deployed and used in healthcare, where doctors and patients tend to trust the models with potent reliability and interpretability. This stringent trustworthiness requirement of the medical field has catalyzed research into Explainable Artificial Intelligence (XAI) for medical image analysis [18, 19, 20], with concept-based methods as one of the representative explainable approaches [20, 21, 22]. These methods integrate human-understandable concepts into the model development and evaluation, offering concept-based explanations and improving model interpretability. Medical concepts can be utilized to make models and their predictions more trustworthy because they provide a meaningful reference for human users by predicting the clinical-relevant attributes and symptoms instead of only giving the diagnosis results. However, most existing generalist medical vision language models ignore the importance of this valuable medical information. Furthermore, fine-grained concept annotations are highly valuable and require human experts’ efforts. Some methods propose to utilize general pre-trained foundation models (e.g., CLIP) to offer concept information [23, 24] but still suffer from low performance in specific domains such as healthcare. Hence building a medical model that is able to annotate concepts in a zero-shot setting holds the potential to facilitate further research for XAI.
@@
In this work, we utilize the fine-grained concept knowledge from the textual data using the UMLS terminology [14]. We evaluate the performance of zero-shot medical concept annotation in various modalities, including dermoscopy, hematology, breast ultrasound, and chest CTs. The results are shown in Table 13. Specifically, our model achieves state-of-the-art performance, outperforming other compared vision-language foundation models. The average performance of ConceptCLIP achieves 10.9% and 10.4% relative performance boost over the results of PMC-CLIP and BiomedCLIP, respectively. Additionally, our proposed model outperforms the domain-specific foundation model MONET on the dermoscopic image datasets. In Fig. 2, we present the image-concept visualization to help better understand the model decision by highlighting the focus region of our model for specific concepts. From the experimental results, it can be observed that our model has the capability to identify clinical concepts across various modalities, thereby holding the potential to facilitate the development of trustworthy models and fine-grained datasets.
Concept-based data analysis
To facilitate trustworthy AI, the quality of the data used is crucial, and the correctness of the correlations within the data impacts model performance. Therefore, it is essential to have a foundation model capable of analyzing correlations between human-understandable concepts and target classes across various image sets. Previous research MONET [3] conducts data auditing to validate the correlations within data. However, it can only be applied to analyze dermoscopic images, which is insufficient for trustworthy AI in general medicine. In contrast, ConceptCLIP can analyze medical images of various modalities by measuring the concept presence difference of different image sets (Method) for medical datasets without fine-grained concept labels. Specifically, to demonstrate the concept-based data analysis ability of our model, we conduct experiments on datasets including modalities of chest X-rays, mammography, pathology, and ultrasound. The results are shown in Fig. 3. We observe that ConceptCLIP successfully identifies differentially present concepts in different datasets. For example, in the analysis for NLM-TB dataset [25], our model concludes that the concepts cavitation, consolidation, opacity, lymphadenopathy, etc. are more likely to present in the image set of tuberculosis while clear lung fields, normal diaphragmatic contours, etc., present more in normal chest X-rays, which are consistent to the human expectation.
Inherently interpretable model employment and evaluation
@@
Within explainable AI, various approaches have been proposed to explain neural networks. Some methods utilize saliency maps [26, 27, 28] to highlight the contribution of each pixel or region in the model’s predictions, while some focus on feature interactions [29], and influence functions [30] to explain the model. However, these post-hoc XAI techniques may not be faithful enough to truly reflect the model decision-making process [31]. Therefore, building inherently interpretable models such as Concept Bottleneck Models (CBMs) [22] gains increasing attention since these models allow doctors and patients to easily understand the model decisions, hence holding the potential to advance trustworthy AI in medicine. Specifically, CBM, as an inherently interpretable model, first predicts the concepts present in the given images and then makes the final prediction based on the concepts through a linear layer. For the concept bottleneck layer, each neuron represents one pre-defined human-understandable concept, thus the weights of the layer can be regarded as the contribution of each concept to the final class (i.e., concept-class association). In this paper, we report the results of inherently interpretable models built upon ConceptCLIP and other vision-language models on disease diagnosis, as shown in Table 14. Our proposed model significantly outperforms all other compared vision-language models, achieving a 6.13% improvement in AUC score over the second-best method. Additionally, the results of concept-class association are shown in Fig. 4, where we present the top 3 relevant concepts for each selected disease across four image modalities. The results of inherently interpretable model employment demonstrate that ConceptCLIP achieves promising diagnosis performance while offering concept-based explanations.
Disease-level concept inspection
@@
In healthcare, concept-based explanations can help medical experts and patients understand the decision process of AI models. In addition to the analysis for a single dataset, a global concept inspection for a disease beyond the granularity of the dataset level is essential for human users to better understand how the model deciphers the disease in terms of human-interpretable concepts. The inspection results may be biased when only one dataset is considered, due to potential spurious correlations within data and the limited number of data samples. The disease-level concept inspection result of our model is shown in Fig. 5, with pneumonia as the selected disease. It can be observed that when only one dataset (e.g., RSNA) is inspected, the output top relevant concepts for pneumonia include consolidation, infiltrates, atelectasis, and effusion. The result changes to the concepts of consolidation, infiltrates, atelectasis, and opacities when more images of pneumonia extracted from different datasets and institutions are considered (e.g., RSNA, VinDr-CXR, VinDr-PCXR). The results of disease-level concept inspection across multiple datasets are more consistent with the ground-truth concepts given by a certified radiologist collaborator. The observations demonstrate the potential of our model to inspect specific diseases with more accurate and consistent global concept-based explanations beyond a single dataset.
@@ Discussion
The introduction of ConceptCLIP, a medical vision-language pre-training model, marks a significant advancement in the field of medical AI, particularly in the realms of image analysis and explainability. By integrating image-text and patch-concept alignments, ConceptCLIP has demonstrated superior performance across a diverse range of medical imaging tasks. However, alongside these accomplishments, the experimental results have also highlighted certain limitations that provide valuable insights for future research and development.
One notable strength of ConceptCLIP is its ability to perform zero-shot medical image diagnosis and cross-modal retrieval with high accuracy. This capability underscores the model’s robustness and adaptability across various medical image modalities without requiring extensive task-specific fine-tuning. However, despite these strengths, the model’s performance in certain complex tasks, such as whole-slide image analysis and visual question answering, suggests room for improvement. These tasks often involve intricate visual and contextual information that might not be fully captured by the current alignment mechanisms. Enhancing the model’s ability to understand and integrate complex visual patterns and contextual cues could further improve its effectiveness in these areas.
Moreover, while ConceptCLIP excels in leveraging UMLS concepts for enhanced interpretability, the reliance on predefined medical concepts may limit the model’s flexibility in adapting to novel or rare medical conditions that are not well-represented in existing medical ontologies. Future iterations of the model could benefit from incorporating more dynamic and adaptive learning strategies that allow for the integration of new concepts as they emerge in medical literature and practice. This could be achieved by employing continual learning frameworks or by integrating external knowledge bases that are frequently updated.
Another limitation arises from the data used for pre-training ConceptCLIP. Although the MedConcept-23M dataset provides a comprehensive resource of medical image-text pairs, it is derived from a specific subset of scientific articles, which may not fully capture the diversity and variability of real-world clinical data. This constraint could affect the generalizability of the model to different clinical settings or populations. Expanding the dataset to include more diverse sources of medical images and texts, such as electronic health records or data from different geographical regions, could enhance the model’s applicability and robustness.
In terms of explainability, while ConceptCLIP offers promising capabilities for concept annotation and inherently interpretable model construction, the evaluation of these features is largely dependent on the accuracy and relevance of the UMLS concepts. The current evaluation metrics may not fully capture the nuanced understanding required for clinical decision-making. Developing more sophisticated evaluation frameworks that consider clinical utility and relevance could provide a more comprehensive assessment of the model’s explainability.
In conclusion, ConceptCLIP represents a significant step forward in the integration of analysis and explainability in medical AI. However, addressing the identified limitations through methodological enhancements and broader data integration will be crucial for realizing the full potential of ConceptCLIP in advancing trustworthy AI in medicine. Future research should focus on refining the model’s adaptability, expanding its data sources, and developing more comprehensive evaluation metrics to ensure its effectiveness and reliability in diverse clinical contexts.
@@
In this section, we first provide a detailed description of the construction of medical image-text-concept triples. Subsequently, we explore how these triples can be utilized for model pre-training.
MedConcept-23M: a dataset of 23 million medical image-text-concept triples
In this section, we detail the construction of the MedConcept-23M dataset, which encompasses two primary stages: dataset collection and concept extraction. Following this, we examine the statistical properties of the proposed MedConcept-23M.
@@
To collect large-scale medical image-text pairs for pre-training a robust medical vision-language model, we focus on the PubMed Central Open Access Subset (PMC-OA)[15], which contains 6,246,351 articles under licenses permitting reuse (as of August 19, 2024). The articles are downloaded from https://ftp.ncbi.nlm.nih.gov/pub/pmc/oa_package/. Each article includes original images and parsed XML files. Utilizing this information, we employ PubMed Parser[32] to extract image-text pairs from each article, yielding a total of 23,289,898 pairs. We allocate 23,000,000 pairs for model pre-training and 289,898 pairs for validation and testing. For the latter set, we apply an image classifier [33] and InternVL [34] to rigorously filter out images containing non-medical information. This process results in the formation of PMC-9K, a cross-modal retrieval dataset in the medical domain for evaluation purposes.
@@
To leverage the knowledge embedded in UMLS for enhancing model pre-training, we extract UMLS concepts from each image caption, thereby coupling the image-text pair with its corresponding UMLS concepts. Specifically, for each caption, we employ SciSpacy [35] linked to the UMLS embedding database to extract entities directly associated with specific UMLS concepts. We filter out recognized UMLS concepts with a matching similarity below 0.8, ultimately extracting relevant UMLS concepts for each caption. This process results in the formation of the MedConcept-23M dataset, comprising 23 million medical image-text-concept triples.
@@
As illustrated in Figure 1 (a), the curated dataset encompasses medical images from various domains, including Radiology, Photography, Microscopy, etc. These images form the foundation for pre-training a robust medical vision-language model capable of handling diverse image modalities.
ConceptCLIP: large-scale medical vision-language model with concept-enhanced pre-training
In this section, we first introduce the model architecture employed for pre-training. Building on this model, we then describe the training objectives in detail, including global image-caption alignment and local patch-concept alignment.
Consider a dataset comprising $N$ image-caption-concept triples, denoted as $\mathcal{D}=\{(I_{1},T_{1},G_{1}),(I_{2},T_{2},G_{2}),\ldots,(I_{N},T_{N},G_{N})\}$, where $I_{m}$ represents the $m$-th image, $T_{m}$ is the $m$-th caption, and $G_{m}$ consists of the concepts extracted from the $m$-th caption. Our objective is to develop a medical vision-language model utilizing an image encoder $f(\cdot)$ and a text encoder $g(\cdot)$. Specifically, for a given image-text pair $(I_{m},T_{m})$, the Vision Transformer-based image encoder $f(\cdot)$ and the BERT-based text encoder $g(\cdot)$ are employed to encode the pair as follows:
$\mathbf{i}_{m}=f(I_{m})$ (1)
$\mathbf{t}_{m}=g(T_{m})$ (2)
Here, $\mathbf{i}_{m}=\{\mathbf{i}_{\text{cls}},\mathbf{i}_{1},\mathbf{i}_{2},...,\mathbf{i}_{r}\}$ represents the encoded representation of the $m$-th image, where $\mathbf{i}_{m}\in\mathbb{R}^{(r+1)\times h}$. $\mathbf{t}_{m}=\{\mathbf{t}_{\text{cls}},\mathbf{t}_{1},\mathbf{t}_{2},...,\mathbf{t}_{s}\}$, where $\mathbf{t}_{m}\in\mathbb{R}^{(s+1)\times h}$ represents the encoded representation of the $m$-th text, $r$ is the number of image patches, $s$ is the number of text tokens, $h$ is the hidden size, $\mathbf{i}_{\text{cls}},\mathbf{t}_{\text{cls}}$ are the global representation of an image and a text sequence, respectively.
Training strategy
@@
In this section, we introduce a dual alignment training strategy, i.e., image-text alignment learning and patch-concept alignment learning.
Image-text alignment learning (IT-Align).
@@
We take sigmoid loss following the idea of SigLIP [5] to conduct the image-text alignment:
$\mathcal{L}_{\text{IT-Align}}=-\frac{1}{|\mathcal{B}|}\sum^{|\mathcal{B}|}_{i=1}\sum^{|\mathcal{B}|}_{j=1}\text{log}\frac{1}{1+e^{{z^{ij}(-t\mathbf{x}_{i}\cdot\mathbf{y}_{j}+b)}}}$ (3)
@@
where $|\mathcal{B}|$ is the size of the mini-batch of image-text pairs, $t$ and $b$ are the logit scale and logit bias for the image-text alignment. $z_{ij}=1$ if the image and text are pairs, and $z_{ij}=-1$ otherwise. $\mathbf{x}_{i}=\frac{\mathbf{i}_{i,\text{cls}}}{||\mathbf{i}_{i,\text{cls}}||}$ and $\mathbf{y}_{j}=\frac{\mathbf{t}_{j,\text{cls}}}{||\mathbf{t}_{j,\text{cls}}||}$ are the $i$-th and the $j$-th normalized image and text representations respectively. These representation vectors are obtained following the attentive pooling in [5] and mean pooling in [36]. The two vectors are then used to calculate the similarity between them.
Patch-concept alignment learning (PC-Align).
To conduct fine-grained alignment between the image and text, the matching between image patches and medical concepts are explored. As shown in Figure 1 (d), image patches are encoded via the image encoder into $\mathbf{i}_{m}$:
$\mathbf{i}_{m}=f(I_{m})$ (4)
For each text $T_{m}=\{t_{1},t_{2},...,t_{N}\}$, where there are $N$ tokens in $T_{m}$, the text encoder encodes the text into $\mathbf{t}_{m}=\{\mathbf{t}_{m,1},\mathbf{t}_{m,2},...,\mathbf{t}_{m,s}\}$:
$\mathbf{t}_{m}=g(T_{m})$ (5)
Considering there is the $b$-th concept in the $m$-th text is $G_{m,b}$, which has the in-text indices: $\{v_{1},v_{2},...,v_{U}\}$, where the length of the concept is $U$, then through the encoded text $\mathbf{t}_{m}$, the concept representation can be derived:
$\mathbf{g}_{m,b}=\text{mean\_pooling}(\mathbf{t}_{m,v_{1}},\mathbf{t}_{m,v_{2}},...,\mathbf{t}_{m,v_{U}})$ (6)

### v2

@@ Results
In this section, we assess the explainability of ConceptCLIP. As the first explainable biomedical foundation model, ConceptCLIP delivers robust and generalizable explainable AI (XAI) performance, consistently providing clinically meaningful concept-level explanations across diverse medical imaging modalities and diagnostic scenarios. Specifically, we analyze the model explainability from three key perspectives: medical concept annotation, inherently interpretable model employment, and concept-disease correlation discovery.
Figure 4: Explainability experiment results. a, The illustration of zero-shot medical concept annotation. b, The illustration of an inherently interpretable model built upon ConceptCLIP, where a linear concept layer is adopted before the final prediction. c, The comparison results of zero-shot concept annotation on datasets of four modalities, which have fine-grained concept labels, with 95% CI reported. d, The effectiveness of the local alignment strategy, where yellow and gray bars represent the medical concept annotation results with local alignment and without local alignment during inference, respectively. e, Performance comparison of inherently interpretable models built upon medical vision-language models on disease diagnosis tasks, with 95% CI reported. In addition, the orange horizontal line indicated in each sub-figure represents the result of fully fine-tuned black-box models. f, Visualization of zero-shot medical concept annotation. For each example, the focused regions are highlighted based on the gradient of image-concept similarities, where the predicted probabilities of concepts are also presented. g, Examples of concept-class association offered by the inherently interpretable model built upon ConceptCLIP. h, Results of discovery for correlations between concepts and diseases on datasets of various modalities. According to the clinical fact, concepts located above the purple horizontal line have stronger associations with the category indicated on the upper right of each figure (e.g., Tuberculosism and Malignant), while concepts below the purple line show greater relevance to the left category (e.g., Normal and Benign). We calculate the concept presence difference to obtain which concepts are more likely to appear in which image set using the zero-shot concept annotation capability of our model. i, Disease-level concept inspection. For a specific disease, disease-level concept inspection conducts global concept-disease association analysis for images from multiple datasets to get more accurate and consistent concept-based explanations.
@@
ConceptCLIP showcases strong ability in medical concept annotation. Concepts are high-level and human-interpretable attributes or symptoms within images, which are essential for doctors or medical experts to diagnose diseases. How to effectively predict concepts for medical images of different modalities is an important and challenging research direction for XAI. In this experiment, we evaluate the performance of zero-shot medical concept annotation in various modalities, including dermatology, pathology, breast ultrasound, and chest CTs. The illustration of medical concept annotation of our model is shown in Fig. 4 (a), where the concept presence outcome is determined by identifying the candidate prompt of concepts with the highest similarity score to the input image. Here, $\mathbf{t}_{c,\text{cls}}$ and $\mathbf{t}_{c,\text{loc}}$ respectively denote the global and local text representations for each candidate prompt of concepts, while $\mathbf{i}_{m,\text{cls}}$ and $\mathbf{i}_{m,\text{loc}}$ signify the global and local representation of the input image, respectively. The results are reported in Fig. 4 (c) and Extended Data Table 13. Specifically, our model achieves state-of-the-art performance, outperforming other compared multimodal foundation models. The average performance of ConceptCLIP achieves 10.9% and 10.4% relative performance boost ($P<0.001$) over the results of PMC-CLIP and BiomedCLIP, respectively, highlighting its superior capability in medical concept annotation. To demonstrate the effectiveness of local alignment for concept annotation, we conduct ablation experiments and report the results in Fig. 4 (d), where our method with local alignment between local image and concept features during inference consistently outperforms baselines without local alignment. In Fig. 4 (f), we present the image-concept visualization to help better understand the model decision by highlighting the focus region of our model for specific concepts. From the experimental results, it can be observed that our model has the capability to identify clinical concepts across various modalities, thereby holding the potential to facilitate the development of trustworthy models and medical image datasets with fine-grained concept annotations.
@@
ConceptCLIP can be employed as an inherently interpretable model. Within explainable AI, various approaches have been proposed to explain neural networks. Some methods utilize saliency maps [19, 20, 21] to highlight the contribution of each pixel or region in the model’s predictions, while some focus on feature interactions [22] and influence functions [23] to explain the model. However, these post-hoc XAI techniques may not be faithful enough to truly reflect the model decision-making process [24]. Therefore, building inherently interpretable models such as Concept Bottleneck Models (CBMs) [9] gains increasing attention since these models allow doctors and patients to easily understand the model decisions, hence holding the potential to advance trustworthy AI in medicine. Specifically, CBM, as an inherently interpretable model, first predicts the concepts present in the given images and then makes the final prediction based on the concepts through a linear layer, as shown in Fig. 4 (b). For the concept bottleneck layer, each neuron represents one pre-defined human-understandable concept. Thus, the weights of the layer can be regarded as the contribution of each concept to the final class (i.e., concept-class association). In this paper, we report the results of inherently interpretable models built upon ConceptCLIP and other multimodal foundation models on disease diagnosis, as shown in Fig. 4 (e) and Extended Data Table 14. Our proposed model significantly outperforms all other compared foundation models, achieving a 6.13% improvement ($P<0.001$) in AUC score over the second-best method. Moreover, we compare the results of the employed inherently interpretable models with black-box models, as indicated using the orange line in Fig. 4 (e). Specifically, we observe that black-box models consistently outperform other medical foundation models with concept bottlenecks. However, our model performs comparably and even outperforms fully fine-tuned black-box models in most settings. For example, ConceptCLIP with CBM gains 5.58% performance boost compared to the fully fine-tuned black-box models on the LUNA16 dataset. Additionally, the results of concept-class association are shown in Fig. 4 (g), where we present the top 3 relevant concepts for each selected disease across four image modalities, ranked by their weights in the concept linear layer, paired with an example image of the specific disease from each considered dataset. The results of inherently interpretable model employment demonstrate that ConceptCLIP achieves promising diagnosis performance while offering concept-based explanations, holding the potential to be employed as an inherently interpretable model and be practically used in disease diagnosis with explanations.
ConceptCLIP can discover correlations between concepts and diseases. One of the most promising applications of XAI in medicine is the discovery and interpretation of correlations between medical concepts and diseases. By understanding the relationships between various medical concepts (e.g., symptoms and observations in medical images) and diseases, the AI model can better assist doctors in diagnosis with more reliable and accurate concept-based explanations. In addition, given the reference of which concepts are more relevant to which diseases, medical experts and patients can perform possible disease screening more easily, leading to earlier and more precise diagnoses and treatments. Previous research [25] conducts data auditing to validate the correlations between concepts and diseases within data. However, it can only be applied to analyze skin images, which is insufficient for trustworthy AI in general medicine. In contrast, ConceptCLIP can discover the correlations and analyze medical images of various modalities by measuring the concept presence difference of different image sets (as described in the Method section) for medical datasets without fine-grained concept labels. Specifically, to demonstrate the concept-disease correlation discovery ability of our model, we conduct comparison experiments on datasets including modalities of chest X-rays, mammography, pathology, and ultrasound. The results are shown in Fig. 4 (h). We observe that ConceptCLIP successfully identifies differentially present concepts in different datasets. For example, in the experiment of the NLM-TB dataset [26], our model concludes that the concepts cavitation, consolidation, opacity, lymphadenopathy, etc., are more likely to present in the image set of tuberculosis while clear lung fields, normal diaphragmatic contours, etc., present more in normal chest X-rays, which are consistent with the clinical findings.
@@
In addition to the analysis for a single dataset, a global concept inspection for a disease beyond the granularity of the dataset level is also essential for human users to better comprehend how the model deciphers the disease in terms of human-interpretable concepts. The inspection results may be biased when only one dataset is considered, due to potential spurious correlations within the data and the limited number of data samples. The disease-level concept inspection result of our model is shown in Fig. 4 (i), with pneumonia as the selected disease. It can be observed that when only one dataset (e.g., RSNA from the National Institutes of Health) is inspected, the output top relevant concepts for pneumonia include consolidation, infiltrates, atelectasis, and effusion. The result changes to the concepts of consolidation, infiltrates, atelectasis, and opacities when more images of pneumonia extracted from multiple datasets are considered (e.g., VinDr-CXR from the Hospital 108 and the Hanoi Medical University Hospital, VinDr-PCXR from Phu Tho Obstetric & Pediatric Hospital). The results of disease-level concept inspection across multiple datasets are more consistent with the ground-truth concepts given by certified radiologists. The observations demonstrate the potential of our model to inspect specific diseases with more accurate and consistent global concept-based explanations across various datasets.
@@ Discussion
In this work, we introduce ConceptCLIP, the first explainable biomedical foundation model that excels in both medical image analysis and explainability, marking a significant advancement in trustworthy AI for healthcare. Leveraging MedConcept-23M, the largest pre-training dataset with 23 million image-text-concept triplets, ConceptCLIP is designed to understand and interpret complex medical data across diverse modalities. This extensive dataset, derived from the PubMed Central Open Access Subset and enriched with UMLS concepts, provides a rich foundation that allows ConceptCLIP to capture intricate relationships and semantics within medical data. This capability is critical for tasks such as medical image analysis and concept-driven explainable AI, where understanding nuanced details in medical images is essential.
Our dual alignment strategy enhances precision and interpretability. The innovative dual alignment approach in ConceptCLIP—comprising image-text alignment (IT-Align) and region-concept alignment (RC-Align)—is instrumental in enhancing the model’s performance. IT-Align integrates global image and text representations, ensuring that the model can comprehend the broader context of medical images. Meanwhile, RC-Align employs UMLS concept knowledge to reinforce local region-concept associations, allowing for more precise and nuanced understanding of specific image areas. This dual strategy significantly improves ConceptCLIP’s ability to accurately associate medical images with relevant concepts, thereby enhancing both analytical precision and interpretability. Experimental results demonstrate that this approach allows ConceptCLIP to outperform other models in zero-shot settings across various modalities, highlighting its robust capability to generalize without additional training.
ConceptCLIP outperforms existing multi-modal medical foundation models across a broad benchmark. Evaluated on the most comprehensive benchmark with 52 tasks across 10 modalities, ConceptCLIP achieves state-of-the-art results among medical multi-modal foundation models (Fig. 1 (d) and (e)). This extensive evaluation includes diverse tasks such as medical image diagnosis, cross-modal retrieval, and visual question answering, ensuring a rigorous assessment of the model’s capabilities. ConceptCLIP consistently surpasses other models in key performance metrics, demonstrating its versatility and robustness. For instance, in zero-shot classification tasks (Fig. 2 (b) and Extended Data Table 3), ConceptCLIP achieves significant AUC gains compared to other models, underscoring its ability to handle a wide range of medical imaging tasks without the need for extensive fine-tuning.
Explainability is a core strength of ConceptCLIP, fostering trust in AI. A standout feature of ConceptCLIP is its ability to provide concept-based explanations alongside diagnostic results. This transparency is critical for clinical adoption, as it enhances trust among healthcare professionals and patients. By elucidating the rationale behind AI predictions, ConceptCLIP supports reliable decision-making. Our experiments show that ConceptCLIP excels in medical concept annotation (Fig. 4 (c) and Extended Data Table 13) and can effectively highlight relevant image regions (Fig. 4 (c)), providing interpretable insights that are crucial for developing trustworthy AI systems in medicine.
The scale of ConceptCLIP is pivotal to its success. The large scale of ConceptCLIP, with a vision encoder of 400 million parameters—the largest among all vision encoders in multi-modal medical foundation models—and its pre-training on an extensive dataset with 23 million image-text-concept triplets are key to its exceptional performance. This underscores the importance of large models and datasets in advancing medical AI, suggesting that future improvements will likely stem from further scaling these resources, thereby enhancing AI capabilities and trustworthiness. The comprehensive nature of MedConcept-23M ensures that ConceptCLIP can generalize across a wide array of tasks and modalities, setting a new benchmark for what is achievable in medical AI.
Limitations and future directions. While ConceptCLIP demonstrates significant advancements in medical vision-language integration, several limitations warrant consideration. First, despite the extensive scale of MedConcept-23M, the dataset may still lack representation of certain rare diseases or conditions, potentially limiting the model’s ability to generalize across all medical scenarios. This highlights the need for continuous dataset expansion and diversification to ensure comprehensive coverage of medical knowledge. Second, while ConceptCLIP excels in providing concept-based explanations, it requires pre-defined concepts for analysis. Its open-set capability, essential for biomarker discovery, remains unexplored and is a promising future direction. Finally, while ConceptCLIP has been evaluated on a broad benchmark, more real-world clinical validation is essential to fully assess its utility and impact in diverse healthcare environments. This necessitates collaboration with clinical practitioners to ensure the model meets practical needs and integrates seamlessly into existing workflows.
@@
In this section, we first provide a detailed description of the construction of medical image-text-concept triplets. Subsequently, we explore how these triplets can be utilized for model pre-training.
MedConcept-23M: a dataset of 23 million medical image-text-concept triplets
In this section, we detail the construction of the MedConcept-23M dataset, which encompasses two primary stages: dataset collection and concept extraction. Following this, we examine the curated dataset statistics.
@@
To collect large-scale medical image-text pairs for pre-training a robust medical vision-language model, we focus on the PubMed Central Open Access Subset (PMC-OA) [15], which contains 6,246,351 articles under licenses permitting reuse (as of August 19, 2024). The articles are downloaded from https://ftp.ncbi.nlm.nih.gov/pub/pmc/oa_package/. Using PubMed Parser [27], we extract 23,289,898 image-text pairs, allocating 23,000,000 for pre-training and 289,898 for evaluation.
The pre-training set retains all image types, including both clinical images (e.g., radiology) and scientific figures (e.g., diagrams), as the latter provide valuable structural and semantic signals (e.g., anatomical relationships, biomarker trends) beneficial for representation learning [28]. For evaluation purposes, we establish PMC-9K through a rigorous curation process: first applying automated filtering via an image classifier [29] and InternVL [30], followed by manual verification ($>96\%$ relevance, $\kappa=0.89$) to ensure clinical applicability.
@@
We develop a rigorous UMLS concept extraction pipeline comprising three stages: (1) Initial biomedical entity recognition using SciSpacy [31], whose UMLS-linked named entity recognizer demonstrates state-of-the-art performance (0.84 F1) on the BC5CDR benchmark [32]; (2) Semantic disambiguation via SciSpacy’s UMLS linker, which computes concept similarity scores using a pretrained clinical BERT model, retaining only matches exceeding the empirically optimized threshold of 0.8 (achieving 95% precision on internal validation); and (3) Contextual validation with PubMedBERT [27] to ensure concept-caption coherence. This process yields MedConcept-23M—a dataset of 23 million precisely aligned image-text-concept triplets.
@@
As illustrated in Fig. 1 (a), (b), and (c), the curated dataset encompasses medical images from various domains, including pathology, ultrasound, X-Ray, and other medical image modalities. These images form the foundation for pre-training a robust medical vision-language model capable of handling diverse image modalities. We treat standard X-ray (denoted as “X-Ray”) and mammography (breast-optimized X-ray, denoted as “Mammography”) as distinct modalities due to fundamental differences in image resolution, contrast requirements, and anatomical focus—factors critical for diagnostic applications.
ConceptCLIP: concept-enhanced contrastive language-image pre-training in medicine
In this section, we first introduce the architecture of the image and text encoders in ConceptCLIP, followed by a detailed description of the training objectives, including global image-text alignment (IT-Align) and local region-concept alignment (RC-Align). We consider a dataset of $N$ image-text-concept triplets, $\mathcal{D}=\{(I_{1},T_{1},G_{1}),\ldots,(I_{N},T_{N},G_{N})\}$, where $I_{m}$ denotes the $m$-th medical image, $T_{m}$ is the corresponding caption, and $G_{m}$ is the set of concepts extracted from $T_{m}$. Our goal is to develop a medical vision-language model (VLM) that learns aligned representations of medical images and text through a dual-encoder framework, enabling robust analysis and explainability in healthcare applications.
Image and Text Encoding
Image Encoder. The image encoder $f(\cdot)$ is a Vision Transformer (ViT) that processes an input image $I_{m}$ into a structured representation. Following SigLIP [28], the encoder outputs:
$\mathbf{i}_{m}=f(I_{m})=[\mathbf{i}_{m,\text{cls}};\mathbf{i}_{m,1};\mathbf{i}_{m,2};\ldots;\mathbf{i}_{m,r}]\in\mathbb{R}^{(r+1)\times h},$ (1)
where $\mathbf{i}_{m,\text{cls}}\in\mathbb{R}^{h}$ is the global [CLS] token embedding, $\mathbf{i}_{m,1:r}\in\mathbb{R}^{r\times h}$ are patch-level region features, $r$ is the number of image regions, and $h$ is the hidden dimension.
Text Encoder. The text encoder $g(\cdot)$ is a PubMedBERT model [27] that maps a text $T_{m}$ into tokenized embeddings:
$\mathbf{t}_{m}=g(T_{m})=[\mathbf{t}_{m,\text{cls}};\mathbf{t}_{m,1};\mathbf{t}_{m,2};\ldots;\mathbf{t}_{m,s}]\in\mathbb{R}^{(s+1)\times h},$ (2)
where $\mathbf{t}_{m,\text{cls}}\in\mathbb{R}^{h}$ is the global text [CLS] token, $\mathbf{t}_{m,1:s}\in\mathbb{R}^{s\times h}$ are token-level embeddings, and $s$ is the text length.
The image and text encoders are initialized with pre-trained weights from SigLIP and PubMedBERT, respectively, and fine-tuned during training to align medical semantics across modalities.
Training approach
@@
In this section, we introduce a dual alignment training approach, which consists of two components: image-text alignment learning and region-concept alignment learning (Fig. 1 (d) and (e)).
@@
Image-text alignment learning (IT-Align). We employ a sigmoid loss function following SigLIP [28], to perform image-text alignment:
$\mathcal{L}_{\text{IT-Align}}=-\frac{1}{|\mathcal{B}|}\sum^{|\mathcal{B}|}_{m=1}\sum^{|\mathcal{B}|}_{n=1}\text{log}\frac{1}{1+e^{{z^{m,n}(-t_{g}\mathbf{x}_{m}\cdot\mathbf{y}_{n}+b_{g})}}}.$ (3)
@@
Here, $|\mathcal{B}|$ represents the size of the mini-batch of image-text pairs. The parameters $t_{g}$ and $b_{g}$ are the logit scale and logit bias for the global image-text alignment, respectively. The variable $z_{m,n}$ equals 1 if the image and text are paired (e.g., $m=n$), and -1 otherwise (i.e., $m\neq n$). The normalized representations of the $m$-th image and the $n$-th text are given by $\mathbf{x}_{m}=\frac{\mathbf{i}_{m,\text{cls}}}{||\mathbf{i}_{m,\text{cls}}||}$ and $\mathbf{y}_{n}=\frac{\mathbf{t}_{n,\text{cls}}}{||\mathbf{t}_{n,\text{cls}}||}$, respectively.
Region-concept alignment learning (RC-Align). To achieve fine-grained alignment between images and text, we explore the matching between image regions and medical textual concepts.
Let the $j$-th concept in the $n$-th text be $G_{n,j}$, with the concept token indices in the original text represented by $\{v_{c}\}_{c=1}^{U}$, where $(1\leq v_{c}\leq s)$ and $U$ is the length of the concept, i.e., the number of words within the concept. The corresponding concept representation is then given by:
$\mathbf{g}_{n,j}=\text{mean\_pooling}(\mathbf{t}_{(n,v_{1})},\mathbf{t}_{(n,v_{2})},...,\mathbf{t}_{(n,v_{U})}).$ (4)